OpenStudy (anonymous):

Is f(x) = √(x-3)^2+3 a function? How do I know it's a function and its domain and range?

OpenStudy (agreene):

\[f(x)=\sqrt{(x-3)^2}+3\] is that the question? if so... it can be simplified..

OpenStudy (agreene):

\[\sqrt{x-3}^2+3\] \[\cancel{\sqrt{}}x-3\cancel{^2}+3\] \[x-3+3\] \[x\] y=x is a line.

OpenStudy (zarkon):

\[\sqrt{f^2}=|f|\]

OpenStudy (anonymous):

yes that is the question

OpenStudy (zarkon):

\[|x-3|+3\]

OpenStudy (anonymous):

why is it an absolute value, how did you get it like that? o.o

OpenStudy (zarkon):

if you square a real number it is always non-negative (zero or positive) when you take the principal value of the square root of a non-negative number you get a non-negative number

OpenStudy (anonymous):

yup. What do you mean by the principal value of the square root?

OpenStudy (zarkon):

the square root of 4 is \(\pm2\) the principal value of the square root of 4 is 2.